# Guaranteed confidence-band enclosures for PDE surrogates

### v1

@@ Abstract
We propose a method for obtaining statistically guaranteed confidence bands for functional machine learning techniques: surrogate models which map between function spaces, motivated by the need build reliable PDE emulators. The method constructs nested confidence sets on a low-dimensional representation (an SVD) of the surrogate model’s prediction error, and then maps these sets to the prediction space using set-propagation techniques. The result are conformal-like coverage guaranteed prediction sets for functional surrogate models. We use zonotopes as basis of the set construction, due to their well studied set-propagation and verification properties. The method is model agnostic and can thus be applied to complex Sci-ML models, including Neural Operators, but also in simpler settings. We also elicit a technique to capture the truncation error of the SVD, ensuring the guarantees of the method.
@@ 1 Introduction
One struggles to find an engineering or scientific discipline which has remained untouched from the rapid progression in machine learning (ML) and artificial intelligence (AI). Scientific machine learning (Sci-ML), with tailored architectures for physics-based problems, have in particular driven major advancements. Neural Operators (NOs) (14), neural networks which can learn between function spaces, have received attention due to their efficient surrogacy of partial differential equation (PDE) solvers (1), and have seen application in complex problems including weather modelling (13) and plasma physics (11).
However, the wide application of AI methods continues despite widely-shared concerns, as adumbrated by (4)(6), and many others, about the reliability and soundness of these methods. This is an opinion that we share: that methods for AI reliability are underdeveloped in comparison to its progression and wide application.
Although multiple methods exist for uncertainty analysis in AI, little R&D effort has gone into approaches which can provide quantitative safety guarantees on the predictions of AI systems. Bayesian machine learning methods, such as Bayesian neural networks, Gaussian processes, Monte Carlo drop-out, and deep ensembles, are powerful methods which can equip predictions with distributional uncertainties. However, they do not attempt to provide statistical guarantees. It has also recently been shown by (2) that Bayesian posteriors (even exact solutions) can suffer from False Confidence, meaning that by their nature (additivity) they can produce unsafe results, including in practical/real-world risk calculations, as they demonstrate. Although the advancements in Bayesian machine learning methods has improved AI reliability, for the above reasons alternatives should perhaps be sought for safety-critical systems.
@@
Presented with a pre-trained model $\hat{f}:X\mapsto F$, which maps from Euclidean space $X\in\mathbb{R}^{m}$ to a space of functions $F\in\mathcal{F}$, and some additional (calibration) data unseen by the model $Z=(Z_{1},\dots,Z_{n})$ where $Z_{i}=(X_{i},F_{i})$, construct a prediction set $\mathbb{C}^{\alpha}\subset\mathcal{F}$ (a band of functions) guaranteed to enclose a next unseen observation $F_{n+1}$ with a user prescribed confidence level: $\mathbb{P}(F_{n+1}\in\mathbb{C}^{\alpha})\geq 1-\alpha$. The space $\mathcal{F}$ has been discretised by a prior analyst $F_{1}=[F_{1}(y_{1}),\dots,F_{1}(y_{l})]\in\mathbb{R}^{l}$, but rather finely $l\gg 1$.
@@ 1.1 Summary of the method
    Construct nested confidence regions $\mathcal{Z}^{\alpha}$ using $\mathcal{Z}$ and $p_{Z}$, such that $\mathbb{P}(U_{n+1}\in\mathcal{Z}^{\alpha})\geq 1-\alpha$,
@@
    Bound truncation error by taking the Cartesian product of the confidence regions $\mathcal{Z}^{\alpha}$ and a bounding box $E$ of the data of the truncated dimension, $\mathcal{R}^{\alpha}=\mathcal{Z}^{\alpha}\times E$,
@@ 1.2 Related work
We mention three distinct prior methods for producing guaranteed multivariate prediction sets, however we would not be surprised if there are other methods we have missed, since uncertainty quantification in Sci-ML is quite a timely problem.
Copula-based conformal prediction
(16) suggest a method to combine univariate prediction sets obtained from conformal prediction using copulas, powerful aggregation functions used to decompose multivariate distributions into their marginals and dependencies. This work was further extended by (20) to time series. We note however that using copulas to model dependence in high dimensions is challenging, and their proposition may require more advanced techniques, such as vine-copulas, to be applicable to functional surrogates.
Supremum-based conformal prediction
: a quite straightforward way to combine univariate conformal predictors is by taking the supremum over a collection of non-conformity scores. (9) take this idea forward by proposing that the scores of each dimension can be normalised or modulated by some function $\sigma(t)$:
@@
Elliptical-set conformal prediction
(17) propose a multi-target (multivariate) non-conformity score
@@
where $\hat{\Sigma}$ is sample covariance of the surrogate’s prediction error. This non-conformity has a known analytical sublevel-set
@@ 1.3 Contributions
Our contribution is most similar to the last of the above methods, where we predict a multivariate set equipped with a guaranteed $\alpha$-level frequentist performance. Our method diverges as we do not use a non-conformity scoring function, but construct the prediction sets directly on a processed calibration data set.
  • A compact representation of probabilistic uncertainty based zonotopes,
  • Methods to calibrate this representation from data, forming a probabilistic bound on the unknown training distribution,
  • An application of this technique to functional surrogate models, giving guaranteed field enclosures as predictions,
  • A method to account for the dimension reduction truncation error, as required by the framework.
@@ 2 Conformal prediction and consonant belief functions
In this section we briefly describe inductive conformal prediction, for calibrating guaranteed prediction sets, particularly used for ML models (but not exclusively, as this paper exemplifies), and the related idea of belief functions, which we find to be a useful theory for performing computation using the calibrated sets.
@@
(ICP) is a computationally efficient version of conformal prediction (18) for computing a set of possible predictions $\mathbb{C}^{\alpha}:X\mapsto\{\text{subsets of}\;Y\}$ of an underlying machine learning model $\hat{f}:X\mapsto Y$. The prediction set is equipped with the following probabilistic inequality
@@
Most relevant for this paper is how $\mathbb{C}^{\alpha}$ is constructed, and how inequality 1 is guaranteed. In ICP, one compares the predictions of a pre-trained model $\hat{f}$ to a calibration dataset $Z=(Z_{1},Z_{2},\dots,Z_{n})$ where $Z_{i}=(X_{i},Y_{i})$, which is yet unseen by $\hat{f}$. A non-conformity score $s:X\times Y\mapsto\mathbb{R}$ is used to compare $\hat{f}$’s predictions to $Z_{i}$, with large values of $s$ indicating a large disagreement between the prediction and the ground truth. A common score used in regression is $s(x,y)=|\hat{f}(x)-y|$. When applied to the calibration data $\alpha_{i}=s(X_{i},Y_{i})$ gives an ordering of $Z_{i}$ in terms of how well $\hat{f}$ performs. Moreover, the probability that another (exchangeable) sample $\alpha_{n+1}=s(X_{n+1},Y_{n+1})$ performs at least as well as previous $\alpha_{i}$ scores is
$\mathbb{P}\left(\frac{1}{n}\sum^{n}_{i=1}\mathbb{I}[\alpha_{n+1}\leq\alpha_{i}]\leq\epsilon\right)\leq\epsilon,$ (2)
where $\mathbb{I}$ is the indicator function that $\alpha_{n+1}\leq\alpha_{i}$. Therefore, taking a (conservative) quantile of the sorted scores $q=\alpha_{\lceil(1-\epsilon)n\rceil}$, where $\lceil x\rceil$ denotes the ceiling function, and computing the set of $\hat{f}$’s possible values giving at least this score $\mathbb{C}^{\alpha}=\{y\in Y\mid s(x,y)\leq q(\alpha)\}$ gives a set satisfying 1. Note that the prediction regions form a nested family of sets w.r.t $\alpha$: $\mathbb{C}^{\alpha_{1}}\supseteq\mathbb{C}^{\alpha_{2}}$ for $0\leq\alpha_{1}\leq\alpha_{2}\leq 1$.
@@
Computing the level-set of a complicated $s(x,y)$ function may be challenging, so often we opt for quite simple non-conformity scores with known level-sets [^1] E.g., the level set of $|y-\hat{f}(x)|$ is simply $[\hat{f}(x)\pm q]$.. A challenge which is exacerbated for multivariate problems, somewhat limiting application. We therefore take a different approach, and directly construct $\mathbb{C}^{\alpha}$ sets using a parametric nested family of sets which we fit to the calibration data (12). This allows us to design prediction sets tailored to our particular data, and well-suited for multivariate problems. Using an interpretation of $\mathbb{C}^{\alpha}$ as belief functions, we can perform additional computations (for example linear and non-linear transformations) on $\mathbb{C}^{\alpha}$, while still maintaining the guarantee 1. This framework however comes with its own challenges. In some sense we are doing the reverse of conformal prediction, where our challenge is not to compute the level-set $\mathbb{C}^{\alpha}=\{y\in Y\mid s(x,y)\leq q(\alpha)\}$, rather we begin with $\mathbb{C}^{\alpha}$ and need to find the membership values $\alpha_{i}=\sup\{\alpha\in[0,1]\;\mid\;Z_{i}\in\mathbb{C}^{\alpha}\}$ of the calibration data. Depending on what set-representation is used, computing the membership values can be a costly operation.
@@
(5) make a useful connection between conformal prediction and belief functions (19), a generalisation of the Bayesian theory of probability where one can make set-valued probabilistic statements, such as inequality 1 given by conformal prediction.
Belief functions, also called random-sets or Dempster-Shafer structures, are set-valued random variables whose statistical properties (such their cdf, moments, sample realisations, and probability measure) are set-valued. Belief functions form a bound on a collection of partially unknown random variables, often used in robust risk analysis. In particular, the conformal nested prediction sets $\mathbb{C}^{\alpha}$ can be related a consonant belief function, and under this framework transformations $f:X\mapsto Y$ of the imprecise random variable are quite simply:
@@
3 Zonotopic confidence sets
We require a base set-representation to construct the prediction sets $\mathbb{C}^{\alpha}$. For this, we chose a class of convex sets known as zonotopes, for some favorable properties, such as being closed under linear transformations and under Minkowski sums. They generalise intervals, boxes, hyper-rectangles, and all their rotations, and may be represented on the computer in a compact manner. An $n$-dimensional zonotope is completely characterised by a vector in $\mathbb{R}^{n}$ (centre), and a collection of $p\in\mathbb{N}$ vectors in $\mathbb{R}^{n}$ (generators).
@@ 3 Zonotopic confidence sets
Figure 1: Example zonotopes and their generators.
We will use zonotopes as a basis to construct nested confidence regions, giving the probabilistic guarantee  1, calibrated with respect to a particular dataset. For this, we specify a parametric family of nested zonotopes, parameterised by a value $\alpha\in[0,1]$. We find that the following parameterisation is simple, and yields nested sets.
@@
A detailed proof that $\mathcal{Z}^{\alpha}_{p_{Z}}$ forms a nested collection of sets in found in appendix A.2. A quick sketch of the proof is that we show that the half-spaces $H=\{x\in\mathbb{R}^{p}|a{\top}x\leq b\}$ composing the box $[-1,1]^{p}$ are nested $H^{\alpha_{1}}\supseteq H^{\alpha_{2}}$ for $\alpha_{1}\leq\alpha_{2}$ under the transformation $c_{\mathcal{Z}}(1-\alpha)+p_{Z}\alpha+G_{\mathcal{Z}}(1-\alpha)X$.
We also note that $\mathcal{Z}^{\alpha=0}_{p_{Z}}=\mathcal{Z}$, which can be called the range (given it is the largest set in the family), and $\mathcal{Z}^{\alpha=1}_{p_{Z}}=\{p_{Z}\}$ (called the core). Several examples of parametric nested zonotopes are shown in Figure 2.
@@
Figure 3: Three examples of the same nested zonotope family $\mathcal{Z}^{\alpha}_{p_{Z}}$, but with different $s(\alpha)$ functions. The central column shows the projection of the set onto the $X_{1}$ dimension.
We note that inequality 1 can be guaranteed irrespective of the chosen $\mathcal{Z}$ and $p_{Z}$, however the ‘quality’ of the confidence sets (their size w.r.t their $\alpha$ confidence-level) depends on how well $\mathcal{Z}^{\alpha}_{p_{Z}}$ captures $X$’s distribution shape. We therefore require the shape of the $\mathcal{Z}^{\alpha}_{p_{Z}}$ sets be fitted with respect to the dataset. That is given, some data $\{X_{1},\dots,X_{N}\}$ sampled from an unknown distribution $X_{i}\sim F_{X}$, we would like to find an enclosing zonotope such that all $X_{i}\in\mathcal{Z}$. We also require the core $p_{Z}$ to be fitted, which plays the role of defining the region of the highest confidence, the ‘central’ point of the dataset in some sense. That is, all the confidence sets will contract towards this point, and so it is desirable for the region around this point to occupy a high density of samples. This point can be determined in terms of data depth, how deep a point is in a dataset with respect to some metric. This point is somewhat analogous to the Bayesian posterior mode. Methods to fit $\mathcal{Z}^{\alpha}_{p_{Z}}$ are described in Section 3.1.
@@
3.1 Fitting zonotopic confidence sets

### v2

# Guaranteed prediction sets for functional surrogate models

@@ Abstract
We propose a method for obtaining statistically guaranteed prediction sets for functional machine learning methods: surrogate models which map between function spaces, motivated by the need to build reliable PDE emulators. The method constructs nested prediction sets on a low-dimensional representation (an SVD) of the surrogate model’s error, and then maps these sets to the prediction space using set-propagation techniques. This results in prediction sets for functional surrogate models with conformal prediction coverage guarantees. We use zonotopes as basis of the set construction, which allow an exact linear propagation and are closed under Cartesian products, making them well-suited to this high-dimensional problem. The method is model agnostic and can thus be applied to complex Sci-ML models, including Neural Operators, but also in simpler settings. We also introduce a technique to capture the truncation error of the SVD, preserving the guarantees of the method.
@@ 1 Introduction
One struggles to find an engineering or scientific discipline which has remained untouched from the rapid progression in machine learning (ML) and artificial intelligence (AI). Scientific machine learning (Sci-ML), with tailored architectures for physics-based problems, have in particular driven major advancements. Neural Operators (NOs) [Li et al., 2020], neural networks which can learn between function spaces, have received attention due to their efficient surrogacy of partial differential equation (PDE) solvers [Azizzadenesheli et al., 2024], and have seen application in complex problems including weather modelling [Kurth et al., 2023] and plasma physics [Gopakumar et al., 2024b].
However, the wide application of AI methods continues despite widely-shared concerns, as adumbrated by Brundage et al. [2020]Dalrymple et al. [2024], and many others, about the reliability and soundness of these methods. This is an opinion that we share: that methods for AI reliability are underdeveloped in comparison to its progression and wide application.
Although multiple methods exist for uncertainty analysis in AI, little R&D effort has gone into approaches which can provide quantitative safety guarantees on the predictions of AI systems. Probabilistic machine learning methods, such as Bayesian neural networks, Gaussian processes, Monte Carlo drop-out, and deep ensembles, are powerful methods which can equip predictions with distributional uncertainties. These include many notable works in neural PDE surrogates and operator learning, including [Yang et al., 2022, Yang and Perdikaris, 2019, Beltran et al., 2024]
However, they do not attempt to provide statistical guarantees. By statistical guarantee we refer to a provable property of the model’s uncertainty, given some quite weak assumptions about the randomness of the data. In this paper, we pursue a method based on conformal prediction [Vovk et al., 2005, Shafer and Vovk, 2008], which is less committal than purely probabilistic approaches, but can yield such guarantees. Rather than producing a full predictive distribution, the method gives a set-valued prediction $\mathbb{C}^{\alpha}$ equipped with a confidence level $1-\alpha$. This prediction set can be guaranteed to contain the next unobserved true label $Y_{n+1}$, with at least probability $1-\alpha$:
$\mathbb{P}(Y_{n+1}\in\mathbb{C}^{\alpha})\geq 1-\alpha.$
The prediction set $\mathbb{C}^{\alpha}$ is constructed using previously observed data $(X_{i},Y_{i})$, and the guarantee holds if the sequence $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n}),(X_{n+1},Y_{n+1})$ is exchangeable, e.g., if the new data point is drawn from the same joint distribution.
This paper’s goal is to extend this framework to models with function-valued outputs. That is, given an input $X$, we want to predict not a single scalar, but a full function $F$ (large vector or tensor e.g., a time series, a spatial field, or a parameterized curve). Our goal is to construct a set of functions that is guaranteed, with a user-defined probability, to contain the true function associated with a new input. Our primary motivation for this is to develop reliable uncertainty quantification for PDE surrogate models, where the functional data represents solutions from numerical PDE solvers. However, the method could be faithfully applied to non-physics cases. Unlike scalar predictions, functional data can exhibit complex dependencies across their domain (e.g., neighboring time points or spatial locations), which our method must capture to produce meaningful predictions. Finally, we want this guarantee to hold using only a held-out calibration dataset. The following formal setup further describes this problem.
@@
Given a pre-trained model $\hat{f}:X\mapsto F$, which maps to a space of functions $F\in\mathcal{F}$, and some additional (calibration) data unseen by the model $Z=(Z_{1},\dots,Z_{n})$ where $Z_{i}=(X_{i},F_{i})$, construct a prediction set $\mathbb{C}^{\alpha}\subset\mathcal{F}$ (a set of functions) guaranteed to enclose a next unseen observation $F_{n+1}$ with a user prescribed confidence level: $\mathbb{P}(F_{n+1}\in\mathbb{C}^{\alpha})\geq 1-\alpha$. The space $\mathcal{F}$ has been discretised $F_{1}=[F_{1}(y_{1}),\dots,F_{1}(y_{l})]\in\mathbb{R}^{l}$, but rather finely $l\gg 1$.
@@ 1.1 Summary of the method
    Construct nested prediction regions $\mathcal{Z}^{\alpha}$ using $\mathcal{Z}$ and $p_{Z}$, such that $\mathbb{P}(U_{n+1}\in\mathcal{Z}^{\alpha})\geq 1-\alpha$,
@@
    Bound truncation error by taking the Cartesian product of the prediction regions $\mathcal{Z}^{\alpha}$ and a bounding box $E$ of the data of the truncated dimension, $\mathcal{R}^{\alpha}=\mathcal{Z}^{\alpha}\times E$,
@@ 1.2 Related work
While we review several prior methods, our coverage may not be exhaustive due to the rapidly evolving literature, since uncertainty quantification in Sci-ML is quite a timely problem.
Copula-based conformal prediction:
Messoudi et al. [2021] suggest a method to combine univariate prediction sets obtained from conformal prediction using copulas, powerful aggregation functions used to decompose multivariate distributions into their marginals and dependencies. This work was further extended by Sun and Yu [2023] to time series. We note however that using copulas to model dependence in high dimensions is challenging, and their proposition may require advanced copula methods, such as vine-copulas, to be applicable to functional surrogates.
Supremum-based conformal prediction:
a quite straightforward way to combine univariate conformal predictors is by taking the supremum over a collection of non-conformity scores. Diquigiovanni et al. [2022] take this idea forward by proposing that the scores of each dimension can be normalised or modulated by some function $\sigma(t)$:
@@
Quantile-functional conformal prediction:
Ma et al. [2024] propose a quantile neural operator, which they calibrate to give a PAC-style bound on the percentage of points in the function domain that falls within a predicted functional uncertainty set. Our methods diverge as we aim to guarantee the entire function $\mathbb{P}(F_{n+1}\in\mathbb{C}^{\alpha})\geq 1-\alpha$ for all values of $F_{n+1}$ simultaneously. While they control the proportion of the function domain covered (with respect to a uniform sampling of $F_{n+1}$’s domain), we aim for full-function coverage. Our method also differs as it does not require a quantile function (an additional neural operator, with additional training data) to be trained (but we do require SVDs).
Elliptical-set conformal prediction:
Messoudi et al. [2022] propose a multi-target (multivariate) non-conformity score
@@
where $\hat{\Sigma}$ is sample covariance of the surrogate’s prediction error. This non-conformity score has a known analytical sublevel-set
@@ 1.3 Contributions
Our contribution is most similar to the last of the above methods, where we predict a multivariate set equipped with a guaranteed $\alpha$-level frequentist performance. Our method diverges as we do not rely on a non-conformity score; instead, we directly construct prediction sets directly on a processed calibration data set. We summarise our contributions:
  • A conformal prediction method based on zonotopes
  • An application of this technique to functional surrogate models, giving multivariate prediction sets for functional data,
  • A method to account for the dimension reduction truncation error, ensuring the guarantees.
@@ 2 Conformal prediction and consonant belief functions
In this section we briefly describe inductive conformal prediction, for calibrating guaranteed prediction sets, particularly used for ML models (but not exclusively), and the related idea of belief functions, which we find to be a useful theory for performing computation using the calibrated sets.
@@
(ICP) is a computationally efficient version of conformal prediction [Papadopoulos et al., 2002] for computing a set of possible predictions $\mathbb{C}^{\alpha}:X\mapsto\{\text{subsets of}\;Y\}$ of an underlying machine learning model $\hat{f}:X\mapsto Y$. The prediction set is equipped with the following probabilistic inequality
@@
Most relevant for this paper is how $\mathbb{C}^{\alpha}$ is constructed, and how inequality 1 is guaranteed. In ICP, one compares the predictions of a pre-trained model $\hat{f}$ to a calibration dataset $Z=(Z_{1},Z_{2},\dots,Z_{n})$ where $Z_{i}=(X_{i},Y_{i})$, which is yet unseen by $\hat{f}$. A non-conformity score $s:X\times Y\mapsto\mathbb{R}$ is used to compare $\hat{f}$’s predictions to $Z_{i}$, with large values of $s$ indicating a large disagreement between the prediction and the ground truth. A common score used in regression is $s(x,y)=|\hat{f}(x)-y|$. When applied to the calibration data, this yields non-conformity scores $\alpha_{i}=s(X_{i},Y_{i})$ for $i=1,\ldots,n$.
The key insight is that these scores form an empirical distribution, and for a new test point, we can compute a p-value based on this empirical distribution. For a candidate prediction $y$ at test input $x$, the p-value is defined as the proportion of calibration scores that are at least as large as the test score:
$p(x,y)=\frac{1}{n}\sum^{n}_{i=1}\mathbb{I}[s(x,y)\leq\alpha_{i}],$ (2)
where $\mathbb{I}$ is the indicator function. Under exchangeability, this p-value has the property that $\mathbb{P}(p(X_{n+1},Y_{n+1})\leq\epsilon)\leq\epsilon$ for any $\epsilon\in[0,1]$.
The prediction set is then constructed by collecting all candidate values $y$ whose p-values exceed the significance level $\alpha$:
$\mathbb{C}^{\alpha}(x)=\{y\in Y\mid p(x,y)>\alpha\}.$
Equivalently, this can be expressed using a quantile of the empirical distribution of non-conformity scores: taking $q=\alpha_{\lceil(1-\alpha)(n+1)\rceil}$ (the $(1-\alpha)$-quantile of the empirical distribution), we have $\mathbb{C}^{\alpha}(x)=\{y\in Y\mid s(x,y)\leq q\}$. Note that the prediction regions form a nested family of sets w.r.t $\alpha$: $\mathbb{C}^{\alpha_{1}}\supseteq\mathbb{C}^{\alpha_{2}}$ for $0\leq\alpha_{1}\leq\alpha_{2}\leq 1$.
@@
Computing the level-set of a complicated $s(x,y)$ function may be challenging, so often we opt for quite simple non-conformity scores with known level-sets, e.g., the level-set of $|y-\hat{f}(x)|$ is simply $[\hat{f}(x)\pm q]$. A challenge which is exacerbated for multivariate problems, somewhat limiting application. We therefore take a different approach, and directly construct $\mathbb{C}^{\alpha}$ sets using a parametric nested family of sets from the calibration data, a method originally suggested by Gupta et al. [2022]. This allows us to design prediction sets tailored to our particular data, and well-suited for multivariate problems. Using an interpretation of $\mathbb{C}^{\alpha}$ as belief functions, we can perform additional computations (for example linear and non-linear transformations) on $\mathbb{C}^{\alpha}$, while still maintaining the guarantee 1. This framework however comes with its own challenges. In some sense we are doing the reverse of conformal prediction, where our challenge is not to compute the level-set $\mathbb{C}^{\alpha}=\{y\in Y\mid s(x,y)\leq q(\alpha)\}$, rather we begin with $\mathbb{C}^{\alpha}$ and need to find the membership values $\alpha_{i}=\sup\{\alpha\in[0,1]\;\mid\;Z_{i}\in\mathbb{C}^{\alpha}\}$ of the calibration data. Depending on what set-representation is used, computing the membership values can be a costly operation.
@@
Cella and Martin [2022] make a useful connection between conformal prediction and belief functions [Shafer, 1976], a generalisation of the Bayesian theory of probability where one can make set-valued probabilistic statements, such as inequality 1 given by conformal prediction.
Belief functions, also called random-sets or Dempster-Shafer structures, are set-valued random variables whose statistical properties (such their cdf, moments, sample realisations, and probability measure) are set-valued. Belief functions form a bound on a collection of partially unknown random variables, often used in robust risk analysis. In particular, the conformal nested prediction sets $\mathbb{C}^{\alpha}$ can be related a consonant belief function [Dubois and Prade, 1990], and under this framework transformations $f:X\mapsto Y$ of the imprecise random variable are quite simply:
@@
3 Zonotope prediction sets
Before proceeding to functional data, we must describe how we will construct our multivariate prediction sets $\mathbb{C}^{\alpha}$. We will use zonotopes—a class of convex sets offering advantageous properties for high-dimensional settings, including closure under linear transformations and Minkowski sums. Zonotopes generalise intervals, boxes, hyper-rectangles, and all their rotations, and may be represented on the computer in a compact manner. An $n$-dimensional zonotope is completely characterised by a vector in $\mathbb{R}^{n}$ (centre), and a collection of $p\in\mathbb{N}$ vectors in $\mathbb{R}^{n}$ (generators).
@@ 3 Zonotope prediction sets
Figure 1: Example zonotopes in blue, and their associated generator matrices immediately below them.
We will use zonotopes as a basis to construct nested prediction regions, giving the probabilistic guarantee  1, calibrated with respect to a particular dataset. For this, we specify a parametric family of nested zonotopes, parameterised by a value $\alpha\in[0,1]$. We find that the following parameterisation is simple, and yields nested sets.
@@
A detailed proof that $\mathcal{Z}^{\alpha}_{p_{Z}}$ forms a nested collection of sets in found in appendix A.2. A quick sketch of the proof is that we show that the half-spaces $H=\{x\in\mathbb{R}^{p}|a{\top}x\leq b\}$ composing the box $[-1,1]^{p}$ are nested $H^{\alpha_{1}}\supseteq H^{\alpha_{2}}$ for $\alpha_{1}\leq\alpha_{2}$ under the transformation $c_{\mathcal{Z}}(1-\alpha)+p_{Z}\alpha+G_{\mathcal{Z}}(1-\alpha)X$. We also note that $\mathcal{Z}^{\alpha=0}_{p_{Z}}=\mathcal{Z}$, (the largest set in the family), and $\mathcal{Z}^{\alpha=1}_{p_{Z}}=\{p_{Z}\}$ Several examples of parametric nested zonotopes are shown in Figure 2.
@@
Why zonotopes
Although Proposition 1 is simple (and its proof not technically deep), the choice of zonotopes is motivated by two key properties:
  • Closure under Cartesian products. The Cartesian product of zonotopes (or with boxes) yields another zonotope. We use this property in step 5 above (section 1.1) when bounding the truncation error of an SVD.
  • Efficiency under linear transformations. Zonotopes can be projected exactly through linear maps via a matrix multiplication. This makes step 6 above very efficient, where we map uncertainty sets back to the prediction space of the surrogate.
Together, these properties make zonotopes a scalable and tractable choice for constructing prediction sets in high-dimensional and function-valued cases.
@@
Figure 3: Three examples of the same nested zonotope family $\mathcal{Z}^{\alpha}_{p_{Z}}$, but with different $s(\alpha)$ functions. The central column shows the projection of the sets onto the $X_{1}$ dimension.
We note that inequality 1 can be guaranteed irrespective of the chosen $\mathcal{Z}$ and $p_{Z}$, however the ‘quality’ of the prediction sets (their size relative to their $\alpha$ confidence-level) depends on how well $\mathcal{Z}^{\alpha}_{p_{Z}}$ captures $X$’s distribution shape. We therefore require the shape of the $\mathcal{Z}^{\alpha}_{p_{Z}}$ sets be fitted with respect to the dataset. That is given, some data $\{X_{1},\dots,X_{N}\}$ sampled from an unknown distribution $X_{i}\sim F_{X}$, we would like to find an enclosing zonotope such that all $X_{i}\in\mathcal{Z}$. We also require the core $p_{Z}$ to be fitted, which plays the role of defining the region of the highest confidence, the ‘central’ point of the dataset in some sense. That is, all the prediction sets will contract towards this point, and so it is desirable for the region around this point to occupy a high density of samples. This point can be determined in terms of data depth, how deep a point is in a dataset with respect to some metric. This point is somewhat analogous to the Bayesian posterior mode. Methods to fit $\mathcal{Z}^{\alpha}_{p_{Z}}$ are described in Section 3.1.
@@
3.1 Fitting zonotope prediction sets